# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

### v1

@@ 7 Conclusion
Generalised rank-constrained approximation in Frobenius norm of a matrix has had a myriad of applications, for example in signal processing, statistical inverse problems and data transmission systems. This finite-dimensional approximation problem has been analysed in [38, 20, 40], for example. In this work we have extended the analysis in the following ways.
@@ 8 Acknowledgements
The research of the authors has been partially funded by the Deutsche Forschungsgemeinschaft (DFG) Project-ID 318763901 – SFB1294.
@@ A.1 Linear Operators
To accommodate the treatment of unbounded operators, we introduce the following concept of a linear map, c.f also [16, Paragraph X.1].
@@
then we say that $T$ is ‘bounded’, and call $\lVert T\rVert$ its ‘operator norm’. Otherwise, $T$ is called ‘unbounded’. As $T$ is linear, it is bounded if and only if it is continuous. If $T$ is bounded, then it has by continuity a unique bounded extension to $\overline{\operatorname{dom}{T}}$. If $T$ is bounded and $\dim\operatorname{ran}{T}<\infty$, then we say that $T$ is ‘finite-rank’ and define $\operatorname{rank}\left({T}\right)\coloneqq\dim\operatorname{ran}{T}$. If also $S:\mathcal{Z}\rightarrow\mathcal{H}$ is a linear operator for some separable Hilbert space $\mathcal{Z}$, then we define $T+S$ on $\operatorname{dom}{T+S}\coloneqq\operatorname{dom}{S}\cap\operatorname{dom}{T}$, and $TS$ on $\operatorname{dom}{TS}\coloneqq S^{-1}(\operatorname{dom}{T})$. If $S$ is finite-rank, then so is $TS$, which is collected below for future reference.
@@
By $\mathcal{B}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$ and $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ we denote the sets of the linear operators $\mathcal{H}\rightarrow\mathcal{K}$ defined on all of $\mathcal{H}$ that are bounded, compact, finite-rank and have rank at most $r$, respectively. We endow these sets with the operator norm and note that all of them except for $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ are linear spaces. We recall that an operator between Hilbert spaces is compact when it is a limit of finite-rank operators in operator norm. When $\mathcal{H}=\mathcal{K}$ we simply write $\mathcal{B}(\mathcal{H})$, $\mathcal{B}_{0}(\mathcal{H})$, $\mathcal{B}_{00}(\mathcal{H})$ and $\mathcal{B}_{00,r}(\mathcal{H})$. For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ we denote its Hilbert space adjoint by $T^{*}\in\mathcal{B}(\mathcal{K},\mathcal{H})$. That is, $T^{*}$ is the unique operator satisfying $\langle Th,k\rangle=\langle h,T^{*}k\rangle$ for $h\in\mathcal{H},k\in\mathcal{K}$. If $T=T^{*}$, then $T$ is ‘self-adjoint’. If $\langle Th,h\rangle\geq 0$ for all $h\in\mathcal{H}$, then $T$ is ‘positive’. If $T$ is positive and self-adjoint, there exists a unique positive self adjoint square root $T^{1/2}$ with $T^{1/2}T^{1/2}=T$. For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ arbitrary, we have that $T^{*}T$ is positive and self adjoint and can define $\lvert T\rvert\coloneqq(T^{*}T)^{1/2}$. We refer to [16, Sections II.2, II.4, VIII.3] for more on the above notions. The following relation between $T$ and $T^{*}$ holds. Note that it implies that $\ker{T}=\operatorname{ran}{T^{*}}^{\perp}$ and $\ker{T^{*}}=\operatorname{ran}{T}^{\perp}$.
@@
This can be found in for example [16, Theorem VI.1.8]. We also recall part of the closed range theorem, c.f. [16, Theorem VI.1.10].
@@
We also recall the definition of Hilbert–Schmidt operators and the Hilbert–Schmidt norm. If $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ satisfies $\sum_{i}\lVert Te_{i}\rVert^{2}<\infty$ for some orthonormal basis (ONB) $(e_{i})_{i}$ of $\mathcal{H}$, then we recall that $T$ is called a ‘Hilbert–Schmidt’ operator and $\lVert T\rVert_{HS}\coloneqq\left(\sum_{i}\lVert Te_{i}\rVert^{2}\right)^{1/2}$ is independent of the choice of $(e_{i})_{i}$. The linear space consisting of all Hilbert–Schmidt operators $L_{2}(\mathcal{H},\mathcal{K})\coloneqq\{T\in\mathcal{B}(\mathcal{H},\mathcal{K}):\ \sum_{i}\lVert Te_{i}\rVert^{2}<\infty\}$ becomes a Hilbert space, with inner product $\langle T,S\rangle_{HS}\coloneqq\sum_{i}\langle Te_{i},Se_{i}\rangle$, which is again independent of the choice of ONB. We call $\langle\cdot,\cdot\rangle_{HS}$ the ‘trace inner product’, and it induces the Hilbert–Schmidt norm $\lVert\cdot\rVert_{HS}$. For $h\in\mathcal{H},k\in\mathcal{K}$, we can interpret the tensor product $k\otimes h\in\mathcal{H}\otimes\mathcal{K}\simeq L_{2}(\mathcal{H},\mathcal{K})$ as the rank-1 map $\tilde{h}\mapsto\langle\tilde{h},h\rangle k$. If $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is the product of two Hilbert–Schmidt operators, then it is called ‘trace-class’. For a trace-class operator $T$ we define its trace $\operatorname{tr}\,\left({T}\right)\coloneqq\sum_{i}\langle Te_{i},f_{i}\rangle$, where $(e_{i})_{i}$ and $(f_{i})_{i}$ are any ONBs of $\mathcal{H}$ and $\mathcal{K}$ respectively. The value of $\operatorname{tr}\,\left({T}\right)$ does not depend on the choice of these ONBs. If $T\in L_{2}(\mathcal{H},\mathcal{K})$, then $\lvert T\rvert^{1/2}$ is trace-class. We recall that a trace-class operator is Hilbert–Schmidt, and a Hilbert–Schmidt operator is compact. Finally, we notice that by the above definitions, $\langle T,S\rangle_{HS}=\operatorname{tr}\,\left({S^{*}T}\right)$ for $S,T\in L_{2}(\mathcal{H},\mathcal{K})$. For a discussion on Hilbert–Schmidt and trace class operators, we refer to [15, Section 3.18].
@@ A.2 Moore–Penrose inverse
Let $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. The Moore–Penrose inverse $T^{\dagger}$ of $T$ is given by a densely defined linear operator $\mathcal{K}\rightarrow\mathcal{H}$ constructed as follows. The continuous linear map $\tilde{T}\coloneqq{\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\ker{T}^{\perp}}}:\ker{T}^{\perp}\rightarrow\mathcal{K}$ is injective and hence invertible on $\operatorname{ran}{T}$. The inverse $\tilde{T}^{-1}$ is thus well-defined on $\operatorname{ran}{T}$. Also, $\tilde{T}^{-1}$ is bounded if and only if $\operatorname{ran}{T}$ is closed. Let $\operatorname{dom}{T^{\dagger}}\coloneqq\operatorname{ran}{T}\oplus\operatorname{ran}{T}^{\perp}$, which is dense in $\overline{\operatorname{ran}{T}}\oplus\operatorname{ran}{T}^{\perp}=\mathcal{K}$. Since every $x\in\mathcal{H}$ has a unique decomposition $x=h+k$ for $h\in\operatorname{ran}{T}$ and $k\in\operatorname{ran}{T}^{\perp}$, we can define the action of $T^{\dagger}$ on $\operatorname{dom}{T^{\dagger}}$ unambiguously via $T^{\dagger}(h+k)\coloneqq\tilde{T}^{-1}h$. So $T^{\dagger}:\mathcal{K}\rightarrow\mathcal{H}$ is linear.
@@
Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces and $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then,
@@
    $\operatorname{ran}{T^{\dagger}}=\ker{T}^{\perp}$,
@@
    $TT^{\dagger}$ $={\left.\kern-1.2ptP_{\overline{\operatorname{ran}{T}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{T^{\dagger}}}},$ (37)
@@
    $T^{\dagger}$ is a closed operator, i.e. it has closed graph,
@@
    $T^{\dagger}$ is bounded if and only if $T$ has closed range,
@@
As $\gamma_{n}\rightarrow 0$, $S$ is indeed compact as the limit of the finite-rank operators $h\mapsto\sum_{n=1}^{N}\gamma_{n}\langle h,e_{n}\rangle e_{n}$ in $B(\mathcal{H},\mathcal{K})$ when $N\rightarrow\infty$. Also, $\operatorname{ran}{S}=\mathcal{D}$ and $\ker{S}=\{0\}$. If $k\in\mathcal{D}$, then $h\coloneqq\sum_{n}\gamma_{n}^{-1}\langle k,e_{n}\rangle e_{n}=Tk\in\mathcal{H}$ and $k=Sh$. Thus $S^{\dagger}:\mathcal{D}=\mathcal{D}\oplus\mathcal{D}^{\perp}\rightarrow\mathcal{H}$ is given by ${T}$.
@@ A.3 Singular Value Decomposition and Eckart–Young Theorem
This truncation is uniquely defined if and only if one of the following conditions holds: $r\geq\dim{\operatorname{ran}{T}}$ or $\sigma_{r}>\sigma_{r+1}$. The first condition expresses the scenario in which $T$ is a finite-rank operator with rank at most $r$, so that $(T)_{r}=T$. The second condition, $\sigma_{r}>\sigma_{r+1}$, is a spectral gap condition, and implies that $T$ is not finite-rank or that $T$ is finite-rank but its rank is at least $r$. Both conditions hold simultaneously if and only if $\operatorname{rank}\left({T}\right)=r$.
@@
$\min\{\lVert T-S\rVert_{HS}:\ S\in\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})\}.$
The SVD of self-adjoint normal operators reduces to $T=\sum_{i\in\mathcal{I}}\lambda_{i}e_{i}\otimes e_{i}$ with real eigenvalues $\lambda_{i}\rightarrow 0$ and orthonormal basis $(e_{i})_{i}$. This follows from the spectral theorem for compact self-adjoint operators, c.f. [16, Theorem II.5.1]. The Hilbert–Schmidt norm can now be calculated in the following way.
@@
Let $T\in L_{2}(\mathcal{H},\mathcal{K})$, then $\lVert T\rVert_{HS}^{2}=\sum_{i}\sigma_{i}(T^{*}T)=\sum_{i}\sigma_{i}(TT^{*})$.
@@
We have $T^{*}T=\sum_{i}\sigma_{i}(T^{*}T)e_{i}\otimes e_{i}$, with ONB $(e_{i})_{i}$ of eigenvectors and real eigenvalues $(\sigma(T^{*}T)_{i})_{i}$. Hence, $\lVert T\rVert_{HS}^{2}=\operatorname{tr}\,\left({T^{*}T}\right)=\sum_{i.j}\langle T^{*}Te_{i},e_{j}\rangle=\sum_{i}\sigma_{i}(T^{*}T)$. The proof for $TT^{*}$ is analogous. ∎
@@ Appendix B Proofs of results
If $T$ is continuous, then $\ker{T}=T^{-1}(\{0\})$ is closed as $\{0\}$ is closed. For the converse, note that $\ker{T}\subset\operatorname{dom}{T}$ closed implies that $\operatorname{dom}{T}/\ker{T}$ is a normed space and the quotient map $Q:\operatorname{dom}{T}\rightarrow\operatorname{dom}{T}/\ker{T}$, $h\mapsto h+\ker{T}$ is continuous, see [16, Section III.4]. Let us define $S:\operatorname{dom}{T}/\ker{T}\rightarrow\mathcal{K}$, $S(h+\ker{T})=Th$ Then $S$ is a well-defined and linear map satisfying $T=S\circ Q$. So $\operatorname{ran}{S}=\operatorname{ran}{T}$ and if $S(h+\ker{T})=0$, then $Th=0$ so that $h\in\ker{T}$ and $S$ is injective. This implies that $\dim{\operatorname{dom}{T}/\ker{T}}=r$. So $S$ is continuous as linear map defined on a finite-dimensional space, c.f. Proposition [16, Propositions III.3.4]. We conclude that $T$ is continuous as a composition of continuous maps. ∎

### v3

@@ 7 Conclusion
Generalised rank-constrained matrix approximation problems in Frobenius norm have had a myriad of applications, for example in signal processing, statistical inverse problems and data transmission systems. These finite-dimensional approximation problems have been analysed in [38, 20, 40], for example. In this work we have extended the analysis in the following ways.
@@ 8 Acknowledgements
The research of the authors has been partially funded by the Deutsche Forschungsgemeinschaft (DFG) Project-ID 318763901 – SFB1294. The authors thank Francesco Carere (Euro-Mediterranean Center on Climate Change) for helpful discussions.
@@ A.1 Linear Operators
To accommodate the treatment of unbounded operators, we introduce the following concept of a linear map, c.f also [15, Paragraph X.1].
@@
then we say that $T$ is ‘bounded’, and call $\lVert T\rVert$ its ‘operator norm’. Otherwise, $T$ is called ‘unbounded’. As $T$ is linear, it is bounded if and only if it is continuous. If $T$ is bounded, then it has by continuity a unique bounded extension to $\overline{\operatorname{dom}{T}}$. If $T$ is bounded and $\dim\operatorname{ran}{T}<\infty$, then we say that $T$ is ‘finite-rank’ and define $\operatorname{rank}\left({T}\right)\coloneqq\dim\operatorname{ran}{T}$. If also $S:\mathcal{Z}\rightarrow\mathcal{H}$ is a linear operator for some separable Hilbert space $\mathcal{Z}$, then we define $T+S$ on $\operatorname{dom}{T+S}\coloneqq\operatorname{dom}{S}\cap\operatorname{dom}{T}$, and $TS$ on $\operatorname{dom}{TS}\coloneqq S^{-1}(\operatorname{dom}{T})$. If $S$ is finite-rank, then so is $TS$. We record this fact below for future reference.
@@
We endow $\mathcal{B}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$ and $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ with the operator norm. We recall that an operator between Hilbert spaces is compact when it is a limit of finite-rank operators in operator norm. When $\mathcal{H}=\mathcal{K}$ we simply write $\mathcal{B}(\mathcal{H})$, $\mathcal{B}_{0}(\mathcal{H})$, $\mathcal{B}_{00}(\mathcal{H})$ and $\mathcal{B}_{00,r}(\mathcal{H})$. For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ we denote its Hilbert space adjoint by $T^{*}\in\mathcal{B}(\mathcal{K},\mathcal{H})$. That is, $T^{*}$ is the unique operator satisfying $\langle Th,k\rangle=\langle h,T^{*}k\rangle$ for $h\in\mathcal{H},k\in\mathcal{K}$. If $T=T^{*}$, then $T$ is ‘self-adjoint’. If $\langle Th,h\rangle\geq 0$ for all $h\in\mathcal{H}$, then $T$ is ‘positive’. If $T$ is positive and self-adjoint, there exists a unique positive self adjoint square root $T^{1/2}$ with $T^{1/2}T^{1/2}=T$. For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ arbitrary, we have that $T^{*}T$ is positive and self adjoint and can define $\lvert T\rvert\coloneqq(T^{*}T)^{1/2}$. We refer to [15, Sections II.2, II.4, VIII.3] for more on the above notions. The following relation between $T$ and $T^{*}$ holds. Note that it implies that $\ker{T}=\operatorname{ran}{T^{*}}^{\perp}$ and $\ker{T^{*}}=\operatorname{ran}{T}^{\perp}$.
@@
This can be found in for example [15, Theorem VI.1.8]. We also recall part of the closed range theorem, c.f. [15, Theorem VI.1.10].
@@
We also recall the definition of Hilbert–Schmidt operators and the Hilbert–Schmidt norm. If $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ satisfies $\sum_{i}\lVert Te_{i}\rVert^{2}<\infty$ for some ONB $(e_{i})_{i}$ of $\mathcal{H}$, then we recall that $T$ is called a ‘Hilbert–Schmidt’ operator and $\lVert T\rVert_{L_{2}}\coloneqq\left(\sum_{i}\lVert Te_{i}\rVert^{2}\right)^{1/2}$ is independent of the choice of $(e_{i})_{i}$. The linear space consisting of all Hilbert–Schmidt operators $L_{2}(\mathcal{H},\mathcal{K})\coloneqq\{T\in\mathcal{B}(\mathcal{H},\mathcal{K}):\ \sum_{i}\lVert Te_{i}\rVert^{2}<\infty\}$ becomes a Hilbert space, with inner product $\langle T,S\rangle_{L_{2}}\coloneqq\sum_{i}\langle Te_{i},Se_{i}\rangle$, which is again independent of the choice of ONB. We call $\langle\cdot,\cdot\rangle_{L_{2}}$ the ‘trace inner product’, and it induces the Hilbert–Schmidt norm $\lVert\cdot\rVert_{L_{2}}$. For $h\in\mathcal{H},k\in\mathcal{K}$, we can interpret the tensor product $k\otimes h\in\mathcal{H}\otimes\mathcal{K}\simeq L_{2}(\mathcal{H},\mathcal{K})$ as the rank-1 map $\tilde{h}\mapsto\langle\tilde{h},h\rangle k$. If $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is the product of two Hilbert–Schmidt operators, then it is called ‘trace-class’. For a trace-class operator $T$ we define its trace $\operatorname{tr}\,\left({T}\right)\coloneqq\sum_{i}\langle Te_{i},f_{i}\rangle$, where $(e_{i})_{i}$ and $(f_{i})_{i}$ are any ONBs of $\mathcal{H}$ and $\mathcal{K}$ respectively. The value of $\operatorname{tr}\,\left({T}\right)$ does not depend on the choice of these ONBs. If $T\in L_{2}(\mathcal{H},\mathcal{K})$, then $\lvert T\rvert^{1/2}$ is trace-class. We recall that a trace-class operator is Hilbert–Schmidt, and a Hilbert–Schmidt operator is compact. Finally, we notice that by the above definitions, $\langle T,S\rangle_{L_{2}}=\operatorname{tr}\,\left({S^{*}T}\right)$ for $S,T\in L_{2}(\mathcal{H},\mathcal{K})$. For a discussion on Hilbert–Schmidt and trace-class operators, we refer to [16, Section 3.18].
@@ A.2 Moore–Penrose inverse
Let $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. The Moore–Penrose inverse $T^{\dagger}$ of $T$ is given by a densely defined linear operator $\mathcal{K}\rightarrow\mathcal{H}$ constructed as follows. The continuous linear map $\tilde{T}\coloneqq{\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\ker{T}^{\perp}}}:\ker{T}^{\perp}\rightarrow\mathcal{K}$ is injective and hence invertible on $\operatorname{ran}{T}$. The inverse $\tilde{T}^{-1}$ is thus well-defined on $\operatorname{ran}{T}$. Also, $\tilde{T}^{-1}$ is bounded if and only if $\operatorname{ran}{T}$ is closed. Let $\operatorname{dom}{T^{\dagger}}\coloneqq\operatorname{ran}{T}\oplus\operatorname{ran}{T}^{\perp}$, which is dense in $\overline{\operatorname{ran}{T}}\oplus\operatorname{ran}{T}^{\perp}=\mathcal{K}$. Since every $x\in\mathcal{K}$ has a unique decomposition $x=h+k$ for $h\in\operatorname{ran}{T}$ and $k\in\operatorname{ran}{T}^{\perp}$, we can define the action of $T^{\dagger}$ on $\operatorname{dom}{T^{\dagger}}$ unambiguously via $T^{\dagger}(h+k)\coloneqq\tilde{T}^{-1}h$. Thus, $T^{\dagger}:\mathcal{K}\rightarrow\mathcal{H}$ is linear.
@@
Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces and $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then the following hold:
@@
    $\operatorname{ran}{T^{\dagger}}=\ker{T}^{\perp}$.
@@
    $TT^{\dagger}$ $={\left.\kern-1.2ptP_{\overline{\operatorname{ran}{T}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{T^{\dagger}}}}.$ (39)
@@
    $T^{\dagger}$ is a closed operator, i.e. it has closed graph.
@@
    $T^{\dagger}$ is bounded if and only if $T$ has closed range.
@@
As $\gamma_{n}\rightarrow 0$, $S$ is indeed compact as the limit of the finite-rank operators $h\mapsto\sum_{n=1}^{N}\gamma_{n}\langle h,e_{n}\rangle e_{n}$ in $B(\mathcal{H},\mathcal{K})$ when $N\rightarrow\infty$. Also, $\operatorname{ran}{S}=\mathcal{D}$ and $\ker{S}=\{0\}$. If $k\in\mathcal{D}$, then $h\coloneqq\sum_{n}\gamma_{n}^{-1}\langle k,e_{n}\rangle e_{n}=Tk\in\mathcal{H}$ and $k=Sh$. Thus $S^{\dagger}:\mathcal{D}=\mathcal{D}\oplus\mathcal{D}^{\perp}\rightarrow\mathcal{H}$ is given by ${T}$. The condition $k\in\mathcal{D}$ is called the “Picard criterion”, see e.g. [18, Theorem 2.8], and under this condition $S^{\dagger}k$ is well-defined.
@@ A.3 Singular Value Decomposition and Eckart–Young Theorem
This truncation is uniquely defined if and only if one of the following conditions holds: $r\geq\dim{\operatorname{ran}{T}}$ or $\sigma_{r}>\sigma_{r+1}$. The first condition expresses the scenario in which $T$ is a finite-rank operator with rank at most $r$, so that $(T)_{r}=T$. The second condition, $\sigma_{r}>\sigma_{r+1}$, implies that $T$ is not finite-rank or that $T$ is finite-rank but its rank is at least $r$. Both conditions hold simultaneously if and only if $\operatorname{rank}\left({T}\right)=r$.
@@
$\min\{\lVert T-S\rVert_{L_{2}}:\ S\in\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})\}.$
The SVD of self-adjoint compact operators reduces to $T=\sum_{i\in\mathcal{I}}\lambda_{i}e_{i}\otimes e_{i}$ with real eigenvalues $\lambda_{i}\rightarrow 0$ and orthonormal basis $(e_{i})_{i}$, c.f. the spectral theorem for self-adjoint and compact operators [15, Theorem II.5.1]. The Hilbert–Schmidt norm can now be calculated in the following way.
@@
Let $T\in L_{2}(\mathcal{H},\mathcal{K})$, then $\lVert T\rVert_{L_{2}}^{2}=\sum_{i}\sigma_{i}(T^{*}T)=\sum_{i}\sigma_{i}(TT^{*})$.
@@
We have $T^{*}T=\sum_{i}\sigma_{i}(T^{*}T)e_{i}\otimes e_{i}$, with ONB $(e_{i})_{i}$ of eigenvectors and real eigenvalues $(\sigma(T^{*}T)_{i})_{i}$. Hence, $\lVert T\rVert_{L_{2}}^{2}=\operatorname{tr}\,\left({T^{*}T}\right)=\sum_{i.j}\langle T^{*}Te_{i},e_{j}\rangle=\sum_{i}\sigma_{i}(T^{*}T)$. The proof for $TT^{*}$ is analogous. ∎
@@ Appendix B Proofs of results
If $T$ is continuous, then $\ker{T}=T^{-1}(\{0\})$ is closed as $\{0\}$ is closed. For the converse, note that $\ker{T}\subset\operatorname{dom}{T}$ closed implies that $\operatorname{dom}{T}/\ker{T}$ is a normed space and the quotient map $Q:\operatorname{dom}{T}\rightarrow\operatorname{dom}{T}/\ker{T}$, $h\mapsto h+\ker{T}$ is continuous, see [15, Section III.4]. Let us define $S:\operatorname{dom}{T}/\ker{T}\rightarrow\mathcal{K}$, $S(h+\ker{T})=Th$. Then $S$ is a well-defined and linear map satisfying $T=S\circ Q$. Thus $\operatorname{ran}{S}=\operatorname{ran}{T}$, and if $S(h+\ker{T})=0$, then $Th=0$, so that $h\in\ker{T}$ and $S$ is injective. This implies that $\dim{\operatorname{dom}{T}/\ker{T}}=r$. Thus $S$ is continuous as a linear map defined on a finite-dimensional space, c.f. [15, Proposition III.3.4]. We conclude that $T$ is continuous as a composition of continuous maps. ∎